OpenStudy (anonymous):

Please help. I will fan everyone that helps and give a medal. Find the value of x: 9 4 9+n x if — = —, then ——— = — n m n m

OpenStudy (anonymous):

do you now whats the formula or any idea how to solve it

OpenStudy (anonymous):

please tell me what math subject this is

geerky42 (geerky42):

HINT: \(\dfrac{9+n}{n}~~~ = ~~~\dfrac{9}{n}+\dfrac{n}{n} ~~~=~~~ \dfrac{9}{n}+1\)

geerky42 (geerky42):

So \(\dfrac{9+n}{n}~~ =~~ \dfrac{4}{m}+1 ~~=~~ \dfrac{4}{m}+\dfrac{m}{m} ~~=~~ \cdots~?\)

OpenStudy (anonymous):

4 _ m

OpenStudy (anonymous):

Before I put 4 as answer but I got it wrong

geerky42 (geerky42):

\(\dfrac{4}{m}+\dfrac{m}{m} ~~=~~ \dfrac{4+m}{m}\), right?

OpenStudy (anonymous):

yes